Find two consecutive integers such that the sum of 8 times the first integer and 6 times the second integer is 48.No book. Thanks so much!
Answer by funmath(2933)

Let the 1st integer=x
The the 2nd integer=x+1
Therefore your problem to solve is:
8x+6(x+1)=48
8x+6x+6=48
14x+6=48
14x+6-6=48-6
14x=42
14x/14=42/14
x=3
The 1st integer=x=3
The sencond integer=x+1=3+1=4